OpenStudy (anonymous):

A small convenience store has determined that weekly demand for chocolate bars has an approximately bell-shaped distribution, with a mean of 55 bars and standard deviation of 6 bars. If the store begins each week with a supply of 77 chocolate bars, approximately what percentage of weeks will supply run out?

OpenStudy (dumbcow):

first determine the z-score z = (77-55)/6 = 22/6 = 3.66 look up in table, 1-P(X<3.66) = percentage of weeks supply runs out

OpenStudy (dumbcow):

should get .1% so essentially they will never run out